# A-FloPS: Accelerating Diffusion Sampling with Adaptive Flow Path Sampler

### v1

@@ Introduction
Diffusion models have achieved remarkable generative performance across diverse modalities, producing high‑fidelity samples that often rival real‑world data (Rombach et al. 2022; Esser et al. 2024; Ramesh et al. 2022). However, their practical deployment is hindered by substantial sampling latency due to the inherently iterative nature of the generation process. Producing a single high‑quality sample typically requires hundreds of network evaluations, making inference computationally expensive and limiting applicability in real‑time or resource‑constrained scenarios (Ho, Jain, and Abbeel 2020; Song et al. 2021).
@@
The second group consists of training‑free methods, which accelerate sampling by improving the numerical integration of the underlying generative ODEs without modifying model parameters. Early work such as DDIM (Song, Meng, and Ermon 2021)—interpretable as a first‑order discretization of the deterministic ODE formulation of diffusion models—showed that this approach outperforms discretizing the stochastic SDE, especially in low‑NFE regimes. More recent high‑order solvers like DPM‑Solver (Lu et al. 2022, 2023) further improve stability and accuracy, though their performance still degrades when the step count becomes extremely small, leaving room for improvement.
Meanwhile, the Flow Matching (FM) framework (Lipman et al. 2023) has emerged as a compelling paradigm for generative modeling. FM learns a continuous‑time velocity field that deterministically transports a simple base distribution to the target data distribution, and often achieves competitive or superior results with fewer discretization steps (Liu, Gong, and Liu 2022; Esser et al. 2024). For Gaussian base distributions, FM can be interpreted as a special case of a diffusion model under a particular noise schedule and a deterministic sampling path (Patel, DeLoye, and Mathias 2024; Gao et al. 2025). However, this correspondence holds only for a specific schedule; a general diffusion model with an arbitrary noise schedule cannot be directly expressed in FM form.
@@
Building on this foundation, we further propose the Adaptive Flow Path Sampler (A‑FloPS), which incorporates an adaptive decomposition of the velocity field to minimize the temporal variability of the residual component. This adaptivity allows high‑order numerical solvers to operate more effectively, resulting in substantial improvements in both stability and sample quality, particularly in the few‑step regime. The code is available at https://github.com/jinc7461/AFloPS.
@@
Modern diffusion models, originally introduced as discrete-time Markov chains, are now predominantly formulated from the perspective of stochastic differential equations (SDEs) (Song et al. 2021). This continuous-time view facilitates theoretical analysis and allows the use of advanced numerical solvers. In this framework, the forward diffusion process gradually perturbs data into noise according to:
@@
These methods improve sampling efficiency without retraining by designing better numerical solvers for the reverse ODE. DDIM (Song, Meng, and Ermon 2021) reformulates sampling as a deterministic ODE and applies first‑order integration. High‑order solvers like DPM‑Solver (Lu et al. 2022) and DPM‑Solver++ (Lu et al. 2023) exploit the semi‑linear structure for higher accuracy in few‑step generation. UniPC (Zhao et al. 2023) further unifies predictor–corrector schemes for diverse noise schedules, reducing the quality gap to full‑trajectory sampling.
@@
Method
Flow Path Sampler (FloPS)
Theorem 1 (Diffusion‑to‑Flow Transformation).
@@
16:   Calculate ${\boldsymbol{x}}_{t_{n+1}}$ as in  (14)
@@
Figure 1: Qualitative comparison of generated images for the “golden retriever” class using different diffusion samplers at varying inference steps. Each row corresponds to a specific sampler, and each column shows results at a different NFE. All methods are initialized with the same random seed for a fair comparison. Our proposed A-FloPS produces sharper details and more coherent object structures, delivering superior visual quality with fewer sampling steps.
@@
With (9), the flow ODE becomes:
${\boldsymbol{x}}_{t_{n+1}}=\mathrm{e}^{\lambda^{(n)}\Delta t}{\boldsymbol{x}}_{t_{n}}+\int_{t_{n}}^{t_{n+1}}\mathrm{e}^{\lambda^{(n)}(t_{n+1}-\tau)}\boldsymbol{h}({\boldsymbol{x}}_{\tau},\tau;\lambda^{(n)})\,\mathrm{d}\tau,$ (13)
@@
The derivative is estimated via backward finite differences:
@@
Step DDIM UniPC DPM-Solver A-FloPS(proposed) FID IS Prec. Rec. FID IS Prec. Rec. FID IS Prec. Rec. FID IS Prec. Rec. 5 38.124 40.570 0.447 0.438 19.165 42.963 0.598 0.431 22.267 41.222 0.571 0.416 6.977 52.704 0.751 0.465 6 22.449 49.972 0.570 0.456 10.165 57.164 0.699 0.472 11.852 55.736 0.681 0.454 4.947 62.512 0.790 0.486 7 14.650 55.273 0.646 0.463 6.698 61.357 0.752 0.493 7.555 60.163 0.739 0.486 4.196 63.818 0.809 0.492 8 10.325 58.539 0.700 0.481 5.320 63.159 0.780 0.504 5.711 62.326 0.772 0.502 3.753 64.857 0.817 0.500 9 7.669 61.023 0.735 0.489 4.828 64.119 0.794 0.514 4.947 63.681 0.787 0.513 3.647 65.077 0.821 0.516 10 6.234 62.205 0.759 0.506 4.638 64.525 0.803 0.516 4.615 64.187 0.799 0.517 3.440 65.663 0.826 0.521
@@
Step DDIM FloPS (proposed) FID IS Prec. Rec. FID IS Prec. Rec. 5 38.12 40.57 0.4465 0.4381 14.9 (–23) 55.06 (+14) 0.6425 (+0.2) 0.458 (+0.02) 6 22.45 49.97 0.5702 0.456 9.721 (–13) 59.06 (+9.1) 0.7057 (+0.14) 0.4814 (+0.025) 7 14.65 55.27 0.6465 0.4631 7.005 (–7.6) 61.89 (+6.6) 0.7469 (+0.1) 0.4869 (+0.024) 8 10.33 58.54 0.6996 0.4813 5.587 (–4.7) 63.36 (+4.8) 0.7713 (+0.072) 0.4964 (+0.015) 9 7.669 61.02 0.7351 0.4889 4.623 (–3) 64.29 (+3.3) 0.7904 (+0.055) 0.5042 (+0.015) 10 6.234 62.2 0.7591 0.5062 4.11 (–2.1) 65.18 (+3) 0.8026 (+0.044) 0.5052 (–0.001)
@@
We compare A-FloPS with representative training-free samplers: DDIM (Song, Meng, and Ermon 2021), DPM-Solver++ (Lu et al. 2022), and UniPC (Zhao et al. 2023). Following (Peebles and Xie 2023), we use a classifier-free guidance scale of $1.5$. For stability, $\lambda^{(n)}$ is constrained to [1,1][-1,1], and $a,b$ in (15) retain terms only up to second order in $\Delta t$.
@@
To answer this, we integrate the adaptive mechanism into a representative FM-based model, Stable Diffusion v3.5 (d=38). As shown in Table 3, it yields consistent gains across most metrics, with the largest improvements at low NFEs. Although CLIP scores for A‑Euler are slightly lower than Euler at $N=7,9$, this is likely due to CLIP’s known saturation and stylistic bias (Yarom et al. 2023). In contrast, IR and GPT-based metrics—better aligned with human judgment—show substantial improvements. The GPT-based evaluation compares two images from the same prompt, selecting the one with better text alignment and perceptual quality; details are in the appendix. Figure 2 further shows that A‑Euler produces sharper textures and more coherent structures than baseline Euler at $\text{NFE}=5$. Additional examples are in the appendix.
These improvements on a native FM model suggest that the adaptive mechanism is not tied to the specifics of our diffusion-to-flow reparameterization. In fact, the mechanism is agnostic to how the FM velocity field is obtained—whether learned directly, distilled from a diffusion model, or analytically derived. It can thus be applied to a broad range of deterministic generative ODEs, such as rectified flows (Liu, Gong, and Liu 2022) and hybrid diffusion–flow systems. While our experiments focus on image generation, the method naturally extends to other modalities such as audio or video synthesis.

### v2

# A-FloPS: Accelerating Diffusion Models via Adaptive Flow Path Sampler

@@ Introduction
Diffusion models have achieved remarkable generative performance across diverse modalities, producing high‑fidelity samples that often rival real‑world data (Rombach et al. 2022; Esser et al. 2024; Ramesh et al. 2022). However, their practical deployment is hindered by substantial sampling latency due to the inherently iterative nature of the generation process. Producing a single high‑quality sample typically requires hundreds of network evaluations, making inference computationally expensive and limiting applicability in real‑time or resource‑constrained scenarios (Ho et al. 2020; Song et al. 2021b).
@@
The second group consists of training‑free methods, which accelerate sampling by improving the numerical integration of the underlying generative ODEs without modifying model parameters. Early work such as DDIM (Song et al. 2021a)—interpretable as a first‑order discretization of the deterministic ODE formulation of diffusion models—showed that this approach outperforms discretizing the stochastic SDE, especially in low‑NFE regimes. More recent high‑order solvers like DPM‑Solver (Lu et al. 2022, 2023) further improve stability and accuracy, though their performance still degrades when the step count becomes extremely small, leaving room for improvement.
Meanwhile, the Flow Matching (FM) framework (Lipman et al. 2023) has emerged as a compelling paradigm for generative modeling. FM learns a continuous‑time velocity field that deterministically transports a simple base distribution to the target data distribution, and often achieves competitive or superior results with fewer discretization steps (Liu et al. 2022; Esser et al. 2024). For Gaussian base distributions, FM can be interpreted as a special case of a diffusion model under a particular noise schedule and a deterministic sampling path (Patel et al. 2024; Gao et al. 2025). However, this correspondence holds only for a specific schedule; a general diffusion model with an arbitrary noise schedule cannot be directly expressed in FM form.
@@
Building on this foundation, we further propose the Adaptive Flow Path Sampler (A‑FloPS), which incorporates an adaptive decomposition of the velocity field to minimize the temporal variability of the residual component. This adaptivity allows high‑order numerical solvers to operate more effectively, resulting in substantial improvements in both stability and sample quality, particularly in the few‑step regime.
@@
Modern diffusion models, originally introduced as discrete-time Markov chains, are now predominantly formulated from the perspective of stochastic differential equations (SDEs) (Song et al. 2021b). This continuous-time view facilitates theoretical analysis and allows the use of advanced numerical solvers. In this framework, the forward diffusion process gradually perturbs data into noise according to:
@@
These methods improve sampling efficiency without retraining by designing better numerical solvers for the reverse ODE. DDIM (Song et al. 2021a) reformulates sampling as a deterministic ODE and applies first‑order integration. High‑order solvers like DPM‑Solver (Lu et al. 2022) and DPM‑Solver++ (Lu et al. 2023) exploit the semi‑linear structure for higher accuracy in few‑step generation. UniPC (Zhao et al. 2023) further unifies predictor–corrector schemes for diverse noise schedules, reducing the quality gap to full‑trajectory sampling.
@@
Method
Flow Path Sampler (FloPS)
Theorem 1 (Diffusion‑to‑Flow Transformation).
@@
16:   Calculate ${\boldsymbol{x}}_{t_{n+1}}$ as in (14) using the finite-difference form of the derivative in (16)
@@
Figure 1: Qualitative comparison of generated images for the “golden retriever” class using different diffusion samplers across varying inference steps. Each row corresponds to a specific sampler (from top to bottom: DDIM, DPM-Solver++, UniPC, FloPS, and A-FloPS), while each column shows results at a different number of function evaluations (NFE = 5, 6, 7, 8, 9, 25). All methods are initialized with the same random seed to ensure a fair comparison. Our proposed A-FloPS generates sharper textures and more coherent object structures, achieving superior visual fidelity with substantially fewer sampling steps.
@@
To achieve a better trade-off between discretization error and the number of function evaluations (NFEs), we employ a high-order integration scheme for the flow ODE. With (9), the flow ODE can be written as:
${\boldsymbol{x}}_{t_{n+1}}$ $=\mathrm{e}^{\lambda^{(n)}\Delta t}{\boldsymbol{x}}_{t_{n}}$
$+\int_{t_{n}}^{t_{n+1}}\mathrm{e}^{\lambda^{(n)}(t_{n+1}-\tau)}\boldsymbol{h}({\boldsymbol{x}}_{\tau},\tau;\lambda^{(n)})\,\mathrm{d}\tau,$ (13)
@@
In practice, instead of computing the time derivative in (14) explicitly, we use its backward finite-difference approximation:
@@
Hence, the update in (14) is implemented using the finite-difference estimate (16) in place of the time derivative.
@@
Step DDIM UniPC DPM-Solver A-FloPS (proposed) FID IS Prec. Rec. FID IS Prec. Rec. FID IS Prec. Rec. FID IS Prec. Rec. 5 38.124 78.897 0.447 0.438 19.165 124.850 0.598 0.431 22.267 115.066 0.571 0.416 6.977 195.176 0.751 0.465 6 22.449 119.865 0.570 0.456 10.165 172.094 0.699 0.472 11.852 162.066 0.681 0.454 4.947 224.613 0.790 0.486 7 14.650 151.406 0.646 0.463 6.698 206.204 0.752 0.493 7.555 196.617 0.739 0.486 4.196 239.762 0.809 0.492 8 10.325 175.733 0.700 0.481 5.320 225.468 0.780 0.504 5.711 216.763 0.772 0.502 3.753 248.932 0.817 0.500 9 7.669 196.445 0.735 0.489 4.828 237.609 0.794 0.514 4.947 231.484 0.787 0.513 3.647 253.016 0.821 0.516 10 6.234 209.926 0.759 0.506 4.638 243.356 0.803 0.516 4.615 239.502 0.799 0.517 3.440 258.486 0.826 0.521
@@
Step DDIM FloPS (proposed) FID IS Prec. Rec. FID IS Prec. Rec. 5 38.12 78.90 0.4465 0.4381 14.9 (–23) 151.44 (+72.5) 0.6425 (+0.2) 0.458 (+0.02) 6 22.45 119.87 0.5702 0.456 9.721 (–13) 181.12 (+61.3) 0.7057 (+0.14) 0.4814 (+0.025) 7 14.65 151.41 0.6465 0.4631 7.005 (–7.6) 204.47 (+53.1) 0.7469 (+0.1) 0.4869 (+0.024) 8 10.33 175.73 0.6996 0.4813 5.587 (–4.7) 219.57 (+43.8) 0.7713 (+0.072) 0.4964 (+0.015) 9 7.669 196.45 0.7351 0.4889 4.623 (–3) 231.76 (+35.3) 0.7904 (+0.055) 0.5042 (+0.015) 10 6.234 209.93 0.7591 0.5062 4.11 (–2.1) 241.02 (+31.1) 0.8026 (+0.044) 0.5052 (–0.001)
@@
We compare A-FloPS with representative training-free samplers: DDIM (Song et al. 2021a), DPM-Solver++ (Lu et al. 2022), and UniPC (Zhao et al. 2023). Following (Peebles and Xie 2023), we use a CFG scale of $1.5$. For stability, $\lambda^{(n)}$ is constrained to [1,1][-1,1], and $a,b$ in (15) retain terms only up to second order in $\Delta t$.
@@
To answer this, we integrate the adaptive mechanism into a representative FM-based model, Stable Diffusion v3.5 (d=38). As shown in Table 3, it yields consistent gains across most metrics, with the largest improvements at low NFEs. Although CLIP scores for A‑Euler are slightly lower than Euler at $N=7,9$, this is likely due to CLIP’s known saturation and stylistic bias (Yarom et al. 2023). In contrast, IR and GPT-based metrics—better aligned with human judgment—show substantial improvements. The GPT-based evaluation compares two images from the same prompt, selecting the one with better text alignment and perceptual quality; details are in the appendix. Figure 2 further shows that A‑Euler produces sharper textures and more coherent structures than baseline Euler at $\text{NFE}=5$.
These improvements on a native FM model suggest that the adaptive mechanism is not tied to the specifics of our diffusion-to-flow reparameterization. In fact, the mechanism is agnostic to how the FM velocity field is obtained—whether learned directly, distilled from a diffusion model, or analytically derived. It can thus be applied to a broad range of deterministic generative ODEs, such as rectified flows (Liu et al. 2022) and hybrid diffusion–flow systems. While our experiments focus on image generation, the method naturally extends to other modalities such as audio or video synthesis.
@@ Conclusion
Ethics Statement
This work focuses on improving the sampling efficiency of pretrained diffusion models. All experiments are conducted on publicly available datasets with appropriate licenses (ImageNet, COCO). No personally identifiable information or sensitive data is used. The proposed method does not introduce new ethical risks beyond those already associated with generative models, such as potential misuse for generating misleading content. We emphasize that our method should be applied responsibly and strictly for research and non-harmful purposes. No human subjects or animals were involved, and no IRB approval was required.
Acknowledgements
This work was supported by NSAF (Grant No. U2230201). All authors are affiliated with the Department of Electronic Engineering, Tsinghua University. In addition, Y. Gu is also affiliated with the Tsinghua Shenzhen International Graduate School and the Beijing National Research Center for Information Science and Technology at Tsinghua University.